OpenStudy (anonymous):

Find an equation of the parabola that has the indicated vertex and whose graph passes through the given point Vertex: ( 2, -1) ; point: ( 4, -3)

OpenStudy (anonymous):

can someone with sense - please help me

hero (hero):

The vertex form of a parabola is \(y = a(x - h)^2 + k\) You are given (h,k) = (2,-1) and (x,y) = (4,-3) Insert the values in place of the variables of the formula to get: \(-3 = a(4 - 2)^2 - 1\) Then solve for \(a\)